dwayne:

The band is selling snacks during lunch. Nachos are $2 each and burgers are $4 each. You want to buy at least 5 items. You want to spend no more than $16 total.

dwayne:

The band is selling snacks during lunch. Nachos are $2 each and burgers are $4 each. You want to buy at least 5 items. You want to spend no more than $16 total. a.Define the variables b. Write a system of inequality c. Give 2 possible solutions Theres the whole question

AngeI:

Hello welcome to QuestionCove cx The variables are 2 and 4 And we have to find how much of each was bought. So the equation would be 2x+4y=16 X being the number of nachos and y being the number of burgers

dwayne:

thank you so much

dwayne:

but how about the 2 possible solutions

AngeI:

So we could try putting numbers for the x and y For example 2 for x and 3 for y 2(2)+4(3)=16 4+12=16 So that one works I want you to try the other one though And by the way it doesn't say you have to get both items. It just has to be at least 5 (so 5 or more) total items

AngeI:

By the way the equality is 2x+5y≥16 Because no more than 16 dollars should be spent c: (The other equation I used was an temporary thing)